Find the area between two curves: y^2 = 2x + 3 ; y = x


1
Expert's answer
2021-05-31T18:33:14-0400
"y^2=2x+3=>x=\dfrac{1}{2}y^2-\dfrac{3}{2}"

"y=x=>y=\dfrac{1}{2}y^2-\dfrac{3}{2}"

"y^2-2y-3=0"

"(y+1)(y-3)=0"

"Point (-1,-1), Point(3,3)"


"Area=\displaystyle\int_{-1}^{3}|y-(\dfrac{1}{2}y^2-\dfrac{3}{2})|dy"

"=\displaystyle\int_{-1}^{3}(y-\dfrac{1}{2}y^2+\dfrac{3}{2})dy"

"=[\dfrac{1}{2}y^2-\dfrac{1}{6}y^3+\dfrac{3}{2}y]\begin{matrix}
 3 \\
 -1
\end{matrix}"

"=\dfrac{9}{2}-\dfrac{27}{6}+\dfrac{9}{2}-\dfrac{1}{2}-\dfrac{1}{6}+\dfrac{3}{2}"


"=\dfrac{16}{3} (units^2)"

Area is "\dfrac{16}{3}" square units.